What is the correct answer for the place of sign of interrogation out of given alternatives?


5 , 7 , 20 , 21 , 80 , 63 , ? , 189.


(A)300 (B)305 (C)310 (D)315 (E)320

Grade:12

3 Answers

vikas askiitian expert
13 years ago

5  ,         7  ,       20  ,      21  ,     80 ,      63  ,        x    ,      189

 

5 ,         7  ,       5x4 ,      7x3 ,   20x4,    21x3 ,    80x4 ,     63x3

 

x = 80x4 = 320

 

approve if u like my ans

bhanuveer danduboyina
13 years ago

E)320

Let

t1=5 , t2=7 , t3=20 , t4=21 , t5=80 , t6=63 , t7=? , t8=189

t1*4=t3                     t2*3=t4

t3*4=t5                     t4*3=t6

t5*4=320                   t6*3=t8

Neer Varshney
13 years ago

320